### v1

@@ A.3 Orchestration
Pipeline Planner: It validates job configurations and generates execution plans by decomposing evaluation jobs into granular evaluation units (user-proxy, dataset, metric, seed combinations). The planner performs compatibility filtering based on component metadata to ensure valid combinations and produces reproducible plan manifests.
Component Registries: It maintains metadata-aware registries for all pluggable components including user-proxy adapters, datasets, metrics, model clients, and judges. Registries support dynamic component discovery and instantiation through decorator-based registration patterns.
Run Controller: It orchestrates end-to-end evaluation runs by managing execution flow, coordinating backend selection (synchronous, asynchronous, or Ray), tracking progress, and handling run lifecycle including resumption of interrupted runs.
Cache Layer: It provides SQLite-based caching for model responses to reduce redundant API calls and improve evaluation efficiency. The cache uses content-based keys (hashing model name, messages, and configuration) and supports namespace isolation and TTL-based expiration.
Observability Layer: It provides structured logging via structlog and optional OpenTelemetry integration for distributed tracing and metrics collection. The layer supports multiple output formats (JSON, text) and configurable log levels for debugging and monitoring.
A.4 Plugin Components
@@
Model Clients: It wraps provider SDKs (OpenAI, LangChain, Azure, etc.) with a unified interface for LLM invocation. Clients automatically capture telemetry (tokens, latency, cost) and support transparent caching through wrapper composition.
User-Proxy Adapters: It adapts diverse agent frameworks and LLM configurations into a standardized AgentSession interface. Adapters normalize differences in message formats, tool usage patterns, and execution models across frameworks.
Datasets: It loads and normalizes evaluation datasets from various sources (HuggingFace, JSONL files, etc.) into standardized episode objects. Loaders support dataset splitting, sampling limits, and provide metadata about available splits and reference statistics.
Metrics: It implements human-likeness measurements including lexical diversity metrics (MATTR, HD-D, Yule’s K) and LLM-as-judge metrics (GTEval, PI, RNR). Metrics declare compatibility requirements through metadata (task types, judge dependencies, reference statistics needs).
A.5 Task Drivers
@@
Default Single-Turn Task Driver: It executes single-turn interactions where the user-proxy receives episode context and generates a single response. This driver is suitable for simple prompt-completion tasks and turn-level utterance evaluation.
Mirror Conversation Task Driver: It orchestrates multi-turn conversational interactions by executing sequences of user-proxy responses within episodic contexts. This driver supports conversation-level evaluations where human-likeness is assessed across extended interactions.
Table 2: Datasets used for the evaluation. All datasets are open-source and include real human user utterances.
Dataset Domain Description Samples Avg Turns
ChatbotArena Zheng et al. (2023) General-purpose, open-domain conversations where users engage with anonymous LLMs on diverse tasks including creative writing, technical questions, reasoning, and casual dialogue. Extracted from winning model interactions. 195 2.5
ClariQ Aliannejadi et al. (2021) Information-seeking dialogues where users pose ambiguous queries and assistants request clarifications to refine search intent. Task-oriented conversations grounded in specific search facets with concise user responses. 200 7.0
OASST1 Köpf et al. (2023) Multilingual instruction-following dialogues covering diverse tasks (coding, Q&A, creative writing, advice). Linear conversation paths extracted from tree-structured human-generated responses. English-only subset used. 200 3.3
QULAC Aliannejadi et al. (2019) Query clarification dialogues in web search contexts. Short user-assistant exchanges covering faceted queries (multiple interpretations) and navigational queries (specific resource seeking). Concise user utterances. 200 2.0
A.6 API & Interface
@@
Runner Facade: It provides a simplified programmatic API for executing evaluation runs from Python code. The facade abstracts orchestration complexity and exposes high-level methods for planning, execution, and result retrieval with progress tracking.
@@
Command-Line Interface (CLI): It exposes framework functionality through a comprehensive CLI supporting workflow commands (plan, dryrun, run), reporting (report json), run management (runs inspect, runs delete), and cache operations (cache stats, cache purge).
@@
Report Generation: It produces structured evaluation summaries with aggregated statistics including means, confidence intervals, and significance testing. The module supports JSON report output with planned support for HTML visualization.
Appendix B Details on Benchmarking datasets
As stated in the Section 5, we use four open datasets. Table 2 provides detailed information about each of the datasets used for this paper.
B.1 Sampling Datasets for Benchmarking
In Table 2, notice that the number of samples we use for evaluating user-proxy agents (approx. 200) does not exhaust all the samples of the underlying open-source datasets. Instead, we perform stratified sampling over the original datasets to extract the final benchmarking datasets.
For each dataset, we first normalize conversations into alternating user-assistant turn sequences, retaining only English dialogues with at least two turns. We then define dataset-specific stratification keys to ensure broad coverage: OASST1 conversations are bucketed by the number of user turns (short, medium, long), ChatbotArena samples are stratified by language, turn count, and whether the dialogue contains multiple user interactions, QULAC entries are grouped by topic type and facet category, and ClariQ dialogues are distributed across topic buckets and clarification-pair counts. Within each stratum, we allocate samples proportionally to the population size while enforcing a minimum-per-stratum threshold to prevent underrepresentation. Using a fixed random seed for reproducibility, we sample up to 200 samples per dataset, yielding a balanced benchmark that spans diverse conversational structures, topical domains, and interaction patterns without oversampling any single mode.
Figure 9: Human-likeness of user-proxy LLMs across four datasets: higher is better for judge metrics; diversity scores are z-scored to humans (0 is best). Judge & assistant fixed to GPT-4o.
Table 3: Experimental Configuration and Dependency Versions
@@ B.1 Sampling Datasets for Benchmarking
Appendix C Auxiliary User-Proxy Evaluations
Similar to Figure 3, where the judge is fixed to Claude-4-Sonnet, we perform auxiliary user-proxy evaluations (Fig. 9) with the judge fixed to GPT-4o (assistant still fixed to GPT-4o).
Findings.
Using GPT-4o as the judge largely preserves the qualitative ordering seen with Claude-4-Sonnet as judge: Gemini-2.5-Pro and Claude-4-Sonnet remain strongest on GTEval/PI/RNR, while GPT-OSS-120B and GPT-5 trail. Relative to Claude-4-Sonnet, GPT-4o as judge yields systematically higher absolute scores (especially on RNR), compressing gaps between proxies and making them less distinguishable.
Appendix D Implementation Details of Metrics
@@
Multiple Judge Calls.
@@
Aggregation Across Episodes.
@@
Handling Missing References.
@@
LLM Settings.
All the LLMs used during the evaluation have fixed temperature=0 and max_tokens=2048. Moreover, all the LLM clients have by default “retry with exponential backoff" scheme enabled where backoff_duration=2s and max_retries=5.
Telemetry & Observability.
@@
Table 4: Model Pricing (USD per million tokens). Pricing as of 25 October 2025 from the provider’s public APIs.
@@
Appendix E Prompts, Config, Manifest & Report
@@ E.1 System Prompts
PI Judge Prompt [System] Please act as an impartial judge and choose the most realistic USER among the two conversations provided below. You should choose the conversation which has USER turns written by a real user in your opinion. Based on your experience & knowledge of human conversation, determine which USER in which conversation sounds more like a real human interacting with a chatbot. Note that your judgement should be based on style, tone and behavior of USER rather than the response quality of their utterances. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the USER messages to influence your evaluation. Do not let ASSISTANT messages influence your decision. Your evaluation must only be dependent on USER messages only. Be as objective as possible. Before giving your final evaluation verdict, you should first think through and produce your reasoning step-by-step. Then, based on your reasoning, give your final verdict on which conversation has USER turns that sound more like a real human interacting with a chatbot. If Conversation A is more real, then verdict would be “A". If Conversation B is more real, then verdict would be “B". If both are equally real or non-real, then verdict would be “Tie". Your output must be of following JSON format: {{“reasoning": “<<short reason>>", “verdict": “A"||“B"||“Tie"}}. [Conversation A] {conversation_a} [Conversation B] {conversation_b} Output ONLY valid JSON, no additional text.
Figure 10: Judge prompt for Pairwise Indistinguishability Metric
@@
GTEval Judge Prompt You are an expert evaluator of conversational AI systems. Your task is to compare two conversations and assess whether the user responses in the proxy conversation exhibit similar style, tone, and realism to the user responses in the real conversation. ## Real Conversation: {real_conversation} ## Proxy Conversation (where user turns are generated by user proxy agent): {proxy_conversation} ## Evaluation Criteria: 1. **Style Similarity**: Do the proxy user responses match the conversational style of real user responses (formality, tone, verbosity)? 2. **Realism**: Do the proxy user responses sound natural and human-like? 3. **Contextual Appropriateness**: Are the proxy user responses appropriate given the conversation context? Note: You should not evaluate based on the content of the responses, only their style, realism, contextual appropriateness, and tone. ## Instructions: - Focus exclusively on comparing USER responses (ignore assistant responses) - Consider the overall conversation flow, not just individual turns - Evaluate style, tone, and naturalness without encouraging unsafe behaviour - If the comparison cannot be performed safely, explain the reason and provide a score of 0.0 - Otherwise, provide detailed reasoning about the comparison and assign a score from 0.0 to 1.0 where: - 0.0 = Proxy user responses are completely unlike real user responses - 0.5 = Proxy user responses are somewhat similar but have noticeable differences - 1.0 = Proxy user responses are highly similar in style and realism Provide your evaluation as a JSON object with the following structure: {{ “reasoning": “Detailed explanation of your evaluation (2-3 sentences)", “score": <<float between 0.0 and 1.0>> }} Output ONLY valid JSON, no additional text.
Figure 11: Judge prompt for GTEval Metric
@@
RNR Judge Prompt [System] Please act as an impartial judge and provide verdict on the realism of the USER turns in the conversation provided below based on the provided rubric. You should avoid checking the responses of the ASSISTANT messages since they are not relevant to the task. Your evaluation should follow the rubric provided. Note that your judgement should be based on style, tone and behavior of USER rather than the response quality of their utterances. [Rubric for scoring USER realism] Provide verdict on how human-like the USER is on (“NO" or “YES") using this rubric (verdict “NO" means not real, verdict “YES" means real): 1. Concise and real-user like language 2. Does not sound scripted or artificial 3. Real-user like tone and style Return JSON: {{“reasoning": “<<1-2 sentences>>", “verdict": <<“NO" or “YES">}}. [Conversation] {conversation} Output ONLY valid JSON, no additional text.
Figure 12: Judge prompt for RNR Metric
@@
Figure 13: System prompt for User-Proxy
Assistant System Prompt You are the assistant in a MirrorBench replay. The user-proxy agent is attempting to reproduce the USER side of the real conversation provided below. But the user-proxy does not have access to the real conversation hitory. Instead, it only has access to the conversation summary. You need to respond as the assistant. But we are providing you with the real conversation history as context, so you can respond consistently same as (or similar to) the original assistant in the real conversation (you may paraphrase lightly for safety). If user-proxy deviates from the original USER turn or the original response would violate policy, reply helpfully using your own knowledge while remaining consistent with the persona demonstrated so far. Always follow ethical content policies. Paraphrase sensitive content instead of quoting it verbatim, and refuse politely if a request is disallowed. Here is the real conversation for context (the USER turns are from the original conversation): {real_conversation} Now, we will provide you with ongoing conversation with the user-proxy. Please respond as the assistant in this conversation.
Figure 14: System prompt for Assistant
Figure 13 and 14 shows system prompts for user-proxy and assistant respectively. Note that once filled, the system prompt would go into chat history as system message. Thus, we do not see the chat history-related placeholder in these system prompts.
E.2 Evaluation Config, Manifest & Report
Figure 15: Job configuration (abridged). YAML that fully specifies a MirrorBench run: backend and concurrency, proxy/assistant clients, dataset binding, judge metrics (with calibration controls), and the mirror-conversation driver.
Figure 16: Plan manifest (JSON, abridged). Fully resolved, replayable spec listing the (proxy, dataset, metric, seed) units, driver params, versions, and config_hash, enabling deterministic re-runs.
Figure 17: Final run report (JSON, abridged). Aggregate metric scores with 95% CIs, calibration controls (HH/$P$), unit grid, and run metadata for the completed evaluation.

### v2

@@ A.1.3. Orchestration
Pipeline Planner
It validates job configurations and generates execution plans by decomposing evaluation jobs into granular evaluation units (user-proxy, dataset, metric, seed combinations). The planner performs compatibility filtering based on component metadata to ensure valid combinations and produces reproducible plan manifests.
Component Registries
It maintains metadata-aware registries for all pluggable components including user-proxy adapters, datasets, metrics, model clients, and judges. Registries support dynamic component discovery and instantiation through decorator-based registration patterns.
Run Controller
It orchestrates end-to-end evaluation runs by managing execution flow, coordinating backend selection (synchronous, asynchronous, or Ray), tracking progress, and handling run lifecycle including resumption of interrupted runs.
Cache Layer
It provides SQLite-based caching for model responses to reduce redundant API calls and improve evaluation efficiency. The cache uses content-based keys (hashing model name, messages, and configuration) and supports namespace isolation and TTL-based expiration.
Observability Layer
It provides structured logging via structlog and optional OpenTelemetry integration for distributed tracing and metrics collection. The layer supports multiple output formats (JSON, text) and configurable log levels for debugging and monitoring.
A.1.4. Plugin Components
@@
Model Clients
It wraps provider SDKs (OpenAI, LangChain, Azure, etc.) with a unified interface for LLM invocation. Clients automatically capture telemetry (tokens, latency, cost) and support transparent caching through wrapper composition.
User-Proxy Adapters
It adapts diverse agent frameworks and LLM configurations into a standardized AgentSession interface. Adapters normalize differences in message formats, tool usage patterns, and execution models across frameworks.
Datasets
It loads and normalizes evaluation datasets from various sources (HuggingFace, JSONL files, etc.) into standardized episode objects. Loaders support dataset splitting, sampling limits, and provide metadata about available splits and reference statistics.
Metrics
It implements human-likeness measurements including lexical diversity metrics (MATTR, HD-D, Yule’s K) and LLM-as-judge metrics (GTEval, PI, RNR). Metrics declare compatibility requirements through metadata (task types, judge dependencies, reference statistics needs).
A.1.5. Task Drivers
@@
Default Single-Turn Task Driver
It executes single-turn interactions where the user-proxy receives episode context and generates a single response. This driver is suitable for simple prompt-completion tasks and turn-level utterance evaluation.
Mirror Conversation Task Driver
It orchestrates multi-turn conversational interactions by executing sequences of user-proxy responses within episodic contexts. This driver supports conversation-level evaluations where human-likeness is assessed across extended interactions.
A.1.6. API & Interface
@@
It provides a simplified programmatic API for executing evaluation runs from Python code. The facade abstracts orchestration complexity and exposes high-level methods for planning, execution, and result retrieval with progress tracking.
Command-Line Interface (CLI)
@@
It exposes framework functionality through a comprehensive CLI supporting workflow commands (plan, dryrun, run), reporting (report json), run management (runs inspect, runs delete), and cache operations (cache stats, cache purge).
Report Generation
@@
It produces structured evaluation summaries with aggregated statistics including means, confidence intervals, and significance testing. The module supports JSON report output with planned support for HTML visualization.
A.2. Execution Flow
Figure 11. MirrorBench execution flow. The framework decomposes an evaluation job into units $\{U_{1},\ldots,U_{p}\}$; each unit $U_{i}$ iterates over episodes $\{e_{1},\ldots,e_{n}\}$ produced by the dataset adapter and executed via task drivers. Metrics are computed per episode and aggregated within each unit with confidence intervals.
This section details the end-to-end execution that instantiates the pipeline $\Psi_{T}$ (as described in §3.1). Figure 11 shows the path from CLI invocation through planning, orchestration, and nested execution to persisted results, compiling high-level specifications into reproducible evaluation artifacts.
CLI Invocation and Planning
An evaluation job is declared via a configuration file (appendix §B.2) that specifies user proxies $\mathcal{P}=\{\theta_{u_{1}},\ldots,\theta_{u_{q}}\}$, datasets $\mathcal{D}=\{D_{1},\ldots,D_{d}\}$, the metric set $M=\{m_{1},\ldots,m_{k}\}$, a fixed assistant model $\theta_{a}$, and run parameters (backend, observability, maximum concurrency, and a seed set $S=\{s_{1},\ldots,s_{r}\}$). The seed set enables repeated trials of the same combination under distinct random initializations, supporting variance estimation and confidence intervals.
Here $\operatorname{Compat}(\cdot)$ returns true iff the components can be jointly executed under their declared capabilities and requirements (e.g. metric prerequisites, model/back-end constraints). Each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$ is executed and evaluated independently, with models, prompts, and driver assignments fixed at plan time. The planner creates & persists a replayable manifest file (appendix §B.2) capturing all requisite metadata, and exposes a dryrun checkpoint for inspection prior to committing computational resources.
Orchestration and Backend Selection
The Runner instantiates a Run Controller that manages execution state, telemetry, and result persistence. The controller initializes a run-specific SQLite database to store task driver outputs and metric values. Based on the run configuration, the system selects an available backend (sync, async, or distributed) and dispatches the list of units $U$ for execution under the chosen strategy. For each unit $U_{i}\in U$, the backend invokes a unit executor that loads the dataset, instantiates the proxy, initializes the metrics, and configures the task driver.
Unit Execution and Episode Extraction
For each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$, the unit executor loads dataset $D_{y}$ and enumerates its constituent episodes. Each episode $e_{j}$ is an immutable specification of a single reference dialogue (§3.1), with dataset layout and fields given by
$D_{y}$ $=\;\{\,e_{1},e_{2},\ldots,e_{n}\,\},$
$e_{j}$ $=\;\bigl(h_{j}^{\mathrm{ref}},\,g_{j},\,\mathrm{meta}_{j}\bigr),$
Conversation Synthesis via Task Drivers
For each episode $e_{j}$, the task driver orchestrates the synthesis of a proxy-assistant interaction by alternating invocations of the user proxy $\theta_{u_{x}}$ (abbrev. $\theta_{u}$) and the assistant model $\theta_{a}$, matching the reference dialogue length $L_{j}=|h_{j}^{\mathrm{ref}}|$. Let $\hat{h}_{j,t}$ denote the synthetic history after $t$ turns, with $\hat{h}_{j,0}=\langle\,\rangle$. The synthetic rollout for episode $e_{j}$ is:
$\textbf{Initialize:}\quad\hat{h}_{j,0}\leftarrow\langle\,\rangle$
$\textbf{For }t=1,\ldots,L_{j}\textbf{:}$
$r_{j}\;=\;\bigl(\,\hat{h}_{j,L_{j}},\;e_{j},\;\mathrm{telemetry}_{j}\bigr),$
Table 3. Experimental Configuration and Dependency Versions
@@ A.2. Execution Flow
Metric Computation
After producing an episode artifact $r_{j}$, metric $m_{z}$ evaluates $r_{j}$ to yield a scalar value $v_{j,m_{z}}$. Depending on the metric type (e.g., lexical, or LLM-judge), the computation may involve reference comparison, pairwise scoring, or external model inference. Formally,
$v_{j,m_{z}}\;=\;m_{z}\!\bigl(r_{j}\bigr),$
Result Aggregation
After processing all episodes $\{e_{1},\ldots,e_{n}\}$ within unit $U_{i}$, the executor aggregates the metric outputs into summary statistics. For metric $m_{z}$, let $V_{U_{i},m_{z}}=\{v_{1,m_{z}},v_{2,m_{z}},\ldots,v_{n,m_{z}}\}$ denote the set of episode-level scores. The unit-level aggregate includes the mean, standard deviation, and $95\%$ confidence interval:
$\bar{v}_{U_{i},m_{z}}\;$ $=\;\frac{1}{n}\sum_{j=1}^{n}v_{j,m_{z}},$
$s_{U_{i},m_{z}}\;$ $=\;\sqrt{\frac{1}{n}\sum_{j=1}^{n}(v_{j,m_{z}}-\bar{v}_{U_{i},m_{z}})^{2}},$
Iteration and Reporting
The backend executes each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$ in $U=\{U_{1},\ldots,U_{p}\}$. Upon completion, the controller updates the last-run pointer for convenient access and consolidates a unified RunSummary that collates, for every $m_{z}$ and unit $U_{i}$, the tuple $\bigl(\bar{v}_{U_{i},m_{z}},\,s_{U_{i},m_{z}},\,\mathrm{CI}_{95}(U_{i},m_{z})\bigr)$, along with cumulative telemetry and execution metadata. This summary serves as the canonical artifact for subsequent JSON/HTML report generation (appendix §B.2) and cross-proxy comparison. The caching layer eliminates repeated LLM calls across units/runs, and the plan manifest enables exact replays of component instantiation and invocation for reproducible results.
A.3. Implementation Details of Metrics
@@
Multiple Judge Calls
@@
Aggregation Across Episodes
@@
Handling Missing References
@@
LLM Settings
All the LLMs used during the evaluation have fixed temperature=0 and max_tokens=2048. Moreover, all the LLM clients have by default “retry with exponential backoff” scheme enabled where backoff_duration=2s and max_retries=5.
Telemetry & Observability
@@
Table 4. Model Pricing (USD per million tokens). Pricing as of 25 October 2025 from the provider’s public APIs.
@@
Appendix B Prompts, Config, Manifest & Report
@@ B.1. System Prompts
PI Judge Prompt [System] Please act as an impartial judge and choose the most realistic USER among the two conversations provided below. You should choose the conversation which has USER turns written by a real user in your opinion. Based on your experience & knowledge of human conversation, determine which USER in which conversation sounds more like a real human interacting with a chatbot. Note that your judgement should be based on style, tone and behavior of USER rather than the response quality of their utterances. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the USER messages to influence your evaluation. Do not let ASSISTANT messages influence your decision. Your evaluation must only be dependent on USER messages only. Be as objective as possible. Before giving your final evaluation verdict, you should first think through and produce your reasoning step-by-step. Then, based on your reasoning, give your final verdict on which conversation has USER turns that sound more like a real human interacting with a chatbot. If Conversation A is more real, then verdict would be “A”. If Conversation B is more real, then verdict would be “B”. If both are equally real or non-real, then verdict would be “Tie”. Your output must be of following JSON format: {“reasoning”: “<<short reason>>”, “verdict”: “A”||“B”||“Tie”}. [Conversation A] {conversation_a} [Conversation B] {conversation_b} Output ONLY valid JSON, no additional text.
Figure 12. Judge prompt for Pairwise Indistinguishability Metric
@@
GTEval Judge Prompt You are an expert evaluator of conversational AI systems. Your task is to compare two conversations and assess whether the user responses in the proxy conversation exhibit similar style, tone, and realism to the user responses in the real conversation. ## Real Conversation: {real_conversation} ## Proxy Conversation (where user turns are generated by user proxy agent): {proxy_conversation} ## Evaluation Criteria: 1. **Style Similarity**: Do the proxy user responses match the conversational style of real user responses (formality, tone, verbosity)? 2. **Realism**: Do the proxy user responses sound natural and human-like? 3. **Contextual Appropriateness**: Are the proxy user responses appropriate given the conversation context? Note: You should not evaluate based on the content of the responses, only their style, realism, contextual appropriateness, and tone. ## Instructions: - Focus exclusively on comparing USER responses (ignore assistant responses) - Consider the overall conversation flow, not just individual turns - Evaluate style, tone, and naturalness without encouraging unsafe behaviour - If the comparison cannot be performed safely, explain the reason and provide a score of 0.0 - Otherwise, provide detailed reasoning about the comparison and assign a score from 0.0 to 1.0 where: - 0.0 = Proxy user responses are completely unlike real user responses - 0.5 = Proxy user responses are somewhat similar but have noticeable differences - 1.0 = Proxy user responses are highly similar in style and realism Provide your evaluation as a JSON object with the following structure: { “reasoning”: “Detailed explanation of your evaluation (2-3 sentences)”, “score”: <<float between 0.0 and 1.0>> } Output ONLY valid JSON, no additional text.
Figure 13. Judge prompt for GTEval Metric
@@
RNR Judge Prompt [System] Please act as an impartial judge and provide verdict on the realism of the USER turns in the conversation provided below based on the provided rubric. You should avoid checking the responses of the ASSISTANT messages since they are not relevant to the task. Your evaluation should follow the rubric provided. Note that your judgement should be based on style, tone and behavior of USER rather than the response quality of their utterances. [Rubric for scoring USER realism] Provide verdict on how human-like the USER is on (“NO” or “YES”) using this rubric (verdict “NO” means not real, verdict “YES” means real): 1. Concise and real-user like language 2. Does not sound scripted or artificial 3. Real-user like tone and style Return JSON: {“reasoning”: “<<1-2 sentences>>”, “verdict”: <<“NO” or “YES”¿}. [Conversation] {conversation} Output ONLY valid JSON, no additional text.
Figure 14. Judge prompt for RNR Metric
@@
Figure 15. System prompt for User-Proxy
Assistant System Prompt You are the assistant in a MirrorBench replay. The user-proxy agent is attempting to reproduce the USER side of the real conversation provided below. But the user-proxy does not have access to the real conversation history. Instead, it only has access to the conversation summary. You need to respond as the assistant. But we are providing you with the real conversation history as context, so you can respond consistently same as (or similar to) the original assistant in the real conversation (you may paraphrase lightly for safety). If user-proxy deviates from the original USER turn or the original response would violate policy, reply helpfully using your own knowledge while remaining consistent with the persona demonstrated so far. Always follow ethical content policies. Paraphrase sensitive content instead of quoting it verbatim, and refuse politely if a request is disallowed. Here is the real conversation for context (the USER turns are from the original conversation): {real_conversation} Now, we will provide you with ongoing conversation with the user-proxy. Please respond as the assistant in this conversation.
Figure 16. System prompt for Assistant
Figure 15 and 16 show system prompts for user-proxy and assistant respectively. Note that once filled, the system prompt would go into chat history as system message. Thus, we do not see the chat history-related placeholder in these system prompts.
B.2. Evaluation Config, Manifest & Report
Figure 17. Plan manifest (JSON, abridged). Fully resolved, replayable spec listing the (proxy, dataset, metric, seed) units, driver params, versions, and config_hash, enabling deterministic re-runs.
Figure 18. Final run report (JSON, abridged). Aggregate metric scores with 95% CIs, calibration controls (HH/$P$), unit grid, and run metadata for the completed evaluation.
Figure 19. Job configuration (abridged). YAML that fully specifies a MirrorBench run: backend and concurrency, proxy/assistant clients, dataset binding, judge metrics (with calibration controls), and the mirror-conversation driver.